OpenStudy (anonymous):

What polynomial function has the roots 11 and 2i?

OpenStudy (jdoe0001):

multiply the roots to get the polynomial keep in mind that complex roots, 2i, aren't all by their lonesome in roots usually, they have a conjugate companion

OpenStudy (anonymous):

isn't 2i the same thing as the square root of -2? @jdoe0001

OpenStudy (jdoe0001):

no, \(\bf i = \sqrt{-1}\)

OpenStudy (campbell_st):

@destinc1215 \[\sqrt{-4} = \sqrt{4i^2} = \pm 2i\]

OpenStudy (anonymous):

so its 2 multiplied times the square root of -1? @jdoe0001

OpenStudy (jdoe0001):

so you have the roots of 11, that is x = 11 => (x-11) = 0 and you have the root of 2i, that is x = 2i AND IT'S CONJUGATE, x = -2i thus (x-2i) = 0 AND (x+2i) =0

OpenStudy (anonymous):

so x=2i and x=-2i and x=11? @jdoe0001

OpenStudy (jdoe0001):

destinyc1215 yes it's, but as pointed out, it's not alone, it has a conjugate component which comes along and it's also one of the roots, as you can see in campbell_st 's line above

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

then i multiply them all together and get my function?

OpenStudy (jdoe0001):

those are the roots, and the factors the will be (x-11)(x-2i)(x+2i) = 0

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf (x-11)(x-2i)(x+2i) = 0\\ \text{keep in mind that }(a-b)(a+b) = (a^2-b^2)\\ (x-11)\left[(x-2i)(x+2i)\right] \implies (x-11)\left[(x^2-(2i)^2)\right]\)

OpenStudy (jdoe0001):

\(\bf (x-11)[(x^2-(4i^2))]\\ \color{blue}{i = \sqrt{-1}\implies i^2 = \sqrt{(-1)^2} \implies -1}\\ (x-11)(x^2-(4(-1))) \)

OpenStudy (jdoe0001):

expand that, and you'd get your polynomial

OpenStudy (jasmineflvs):

f(x) = x^3 - 11x^2 + 4x - 44 <===ANSWER